rectangle is 6 feet long and 12 feet wide. What length of a simalar rectangle whose width is 4 feet?
Found 2 solutions by stanbon, checkley77:
Answer by stanbon(75887)

rectangle is 6 feet long and 12 feet wide. What length of a simalar rectangle whose width is 4 feet?
---
Proportion:
length/4 = 6/12
length = 4(6/12)
length = 2
=================================
Cheers,
Stan H.

Answer by checkley77(12844)

6:12=4:x cross multiply.
6x=48
x=48/6
x=8 ft. for the width.